How do I solve this. I tried cross multiplying, didn't work.
x ab 9 - x + 2 ab 3 = 4
-- --

Answer by rothauserc(4718)

x ab 9 - x + 2 ab 3 = 4
xab9 - x + 6ab = 4
x(ab9-1) + 6ab = 4
x(ab9-1) = 4-6ab
x = 2(2-3ab) / (ab9-1)